OpenStudy (anonymous):

i cant figure out how to write y= 1/6x+12 in standard form using integers

OpenStudy (anonymous):

move x to the other side and multiply everything by 6 y=(1/6)x+12 -1/6x -1/6x 6(y-(1/6)x=12) 6y-x=72